# HAL-Based Plugin Estimation of the Causal Dose-Response Curve

### v3

@@ Abstract
Estimating and obtaining reliable inference for the marginally adjusted causal dose-response curve for continuous treatments without relying on parametric assumptions is a well-known statistical challenge. Parametric models risk introducing significant bias through model misspecification, compromising the accurate representation of the underlying data and dose-response relationship. On the other hand, nonparametric models face difficulties as the dose-response curve is not pathwise differentiable, preventing consistent estimation at standard rates. The Highly Adaptive Lasso (HAL) maximum likelihood estimator offers a promising approach to this issue. In this paper, we introduce a HAL-based plug-in estimator for the causal dose-response curve and assess its empirical performance against other estimators. Through comprehensive simulations, we evaluate the accuracy of the estimation and the quality of the inference, particularly in terms of coverage, using robust standard error estimators. Our results demonstrate the finite-sample effectiveness of the HAL-based estimator, utilizing an undersmoothed and smoothness-adaptive fit for the conditional outcome model. Additionally, the simulations reveal that the HAL-based estimator consistently outperforms existing methods for estimating the causal dose-response curve.
@@ 1 Introduction
A marginally adjusted dose-response curve, often referred to as a "causal" dose-response curve, provides a graphical representation of the population-level probability of a health outcome following a uniform application of a specific dose, or more generally, a dosage rule applied to the target population. Estimation of such a curve where the dose is ordinal or continuous requires accounting for potential confounders, missingness, and other typical complications in data-generating processes. Estimating the marginally adjusted dose-response is a common problem in causal inference. Efficient estimators with valid inference are crucial for a wide range of scientific applications, including epidemiology, medical research, social sciences, and many other fields. For example, in a clinical trial investigating the impact of a continuous drug dosage on blood pressure reduction, the dose-response curve can provide valuable insights into the specific reduction or increase in blood pressure that is attributed to each incremental change in dosage, accounting for confounding factors such as demographic and baseline clinical characteristics.
@@
To estimate the causal dose-response curve, we need to define a parameter under the counterfactual or potential outcomes framework in which the outcome would be observed for subjects in the target population if all of them received a specific treatment level van der Laan and Petersen (2008); Petersen et al. (2006); Rubin (2004). This framework, widely used in causal inference, has predominantly focused on binary interventions Van der Laan et al. (2011); Rubin (2006); Laan and Robins (2003); Abadie and Imbens (2006); Angrist and Imbens (1995); Robins (1986); Rubin (1978). Common approaches for estimating causal effects, such as the average treatment effect (ATE), include matching Abadie et al. (2011), estimating equations (e.g., inverse probability of treatment weighting (IPTW) Austin and Stuart (2015) and augmented IPTW (A-IPTW) Luque-Fernandez et al. (2018)), and substitution estimators (e.g., G-computation Robins (1986) and targeted maximum likelihood estimation (TMLE) Van der Laan et al. (2011)). Additionally, highly adaptive lasso (HAL) van der Laan (2017) and undersmoothed HAL van der Laan et al. (2022) have been developed to provide reliable inference for binary treatment effects. While significant progress has been made in the binary treatment context, many causal inference questions inherently involve continuous or ordinal interventions, such as doses in clinical treatments, for which consistent and efficient estimators remain scarce.
@@
The primary challenge in estimating the causal dose-response curve for continuous treatments in a nonparametric model is that the parameter is not pathwise differentiable Bickel et al. (1993). Consequently, $\sqrt{n}$ estimators are not applicable, leading to reliance on estimators that depend on data density rather than a smooth function of this density.
@@
Considerable efforts have been dedicated to crafting estimators for the causal dose-response curve. The Marginal Structural Model (MSM) framework, originally introduced by Robins et al. (2000), is known for defining causal dose-response curves in parametric models. However, its efficacy is contingent upon the inclusion of the true dose-response curve within the model, alongside its structural causal assumptions. To address potential model misspecification, enhancements to MSM have been proposed using outcome prediction models, data-adaptively estimation of MSM, and projecting the true causal dose-response curve onto a functional space defined by MSM Robins et al. (2000); Neugebauer and van der Laan (2007). Despite these advancements, challenges related to robustness remain. Alternative estimators based on plug-in methods for conditional outcome models have been proposed Hirano and Imbens (2004); Díaz and van der Laan (2013); Schwab et al. (2020); Callaway et al. (2021), but these often struggle with dimensionality issues, restrictive assumptions, and the complexity of valid inference. Techniques such as cross-validated TMLE (chapter 25 in Van der Laan and Rose (2018) and nonparametric causal inference Kennedy et al. (2017) achieve asymptotic normality, facilitating inference, but they are constrained by their dependence on dimension and smoothness. Additionally, their reliance on kernel smoothing introduces complexities in selecting the tuning parameter, particularly given the suboptimal performance of global cross-validation selectors for bandwidth selection.
@@
In this article, we introduce and refine the HAL-based plugin estimator for the marginally adjusted causal dose-response. HAL-MLE is a nonparametric minimum-loss estimator capable of efficiently estimating pathwise differentiable parameters. Importantly, it has been theoretically proven to perform well with pointwise asymptotic normality in settings where the parameter is non-pathwise differentiable van der Laan (2023), such as with continuous treatments. Specifically, the HAL-based plugin estimator converges to the true curve at a dimension-free rate up to log n factors van der Laan (2023).
@@
Through simulation studies, we assess the performance of this HAL-based estimator under various realistic conditions, showing that it provides valid inference with robust confidence intervals for the causal dose-response curve. Our contribution extends the utility of HAL to continuous dose-response problems, which have traditionally been difficult to estimate efficiently due to non-pathwise differentiability. We present the theoretical underpinnings of the HAL-based plug-in estimator and its associated delta-method-based inference, followed by a series of simulations demonstrating its consistent performance and reliability.
@@
In the sections that follow, we define the target parameter and present the HAL-based plug-in estimator with delta-method-based confidence intervals. Then, we present simulation results that highlight the estimator’s asymptotic normality, nominal coverage, and overall reliability for estimating the marginally adjusted dose-response curve in continuous treatment settings.
@@ 2.1 Target Parameter and Statistical Model
Consider an observational study or experiment with $n$ independent and identically distributed observations of a continuous exposure or treatment variable $A$, an outcome $Y$, and covariates $W$, represented as $O=(W,A,Y)$. $O$ follows a true data-generating distribution (DGD) $P_{0}\in{\cal M}^{NP}$, where $\mathcal{M}^{NP}$ is a realistic nonparametric model. The goal is to estimate the marginal causal dose-response curve, which captures the causal relationship between the treatment and outcome. The target parameter $\Psi_{P}(a)$ is defined as:
@@
[…]
Estimating this target parameter is challenging since there is no efficient estimation method with reliable inference. Therefore, we adopt a flexible and realistic statistical model based on càdlàg functions with finite sectional variation norm. This approach allows for robust estimation without making unrealistic assumptions.
@@ 2.2 Flexible and Realistic Statistical Model: Càdlàg Functions
To model the relationship between $A$, $Y$, and $W$ without relying on strong parametric assumptions, we use the class of multivariate càdlàg (right-continuous with left limits) functions with finite sectional variation norms (FSVN) as the foundation of the statistical model. This class includes not only continuous functions but also functions with jumps and quadratic features, providing high flexibility for approximating complex relationships in real-world data while avoiding strict smoothness assumptions.
@@
$||Q||^{*}_{v}=|Q(0)|+\sum_{S\subset\{1,\dots,d\}}\int_{(0(S),1(S)]}|dQ_{S}(u)|.$
Càdlàg functions form a Donsker class, ensuring that empirical processes indexed by this class converge to Gaussian processes. This convergence provides desirable statistical properties such as optimal convergence rates and asymptotic normality Van der Vaart (1996). These properties make càdlàg functions an attractive choice for constructing estimators in complex, high-dimensional data settings. In particular, Highly Adaptive Lasso (HAL) leverages this flexibility to estimate the target function $Q$ robustly, while achieving optimal convergence rates under weak assumptions.
@@ 2.3 Highly Adaptive Lasso (HAL) and Representation of $Q$: From Statistical Model to Practical Estimator
Building on the flexibility of càdlàg functions, the Highly Adaptive Lasso (HAL) is a nonparametric minimum-loss estimator that can flexibly estimate complex, high-dimensional target parameters. In its most nonparametric form, HAL only assumes that the true target function belongs to the class of $d$-variate real-valued càdlàg functions with bounded sectional variation norms. This allows HAL to fit a wide range of functions, including those with discontinuities and non-linear relationships, making it a powerful tool for data-driven estimation without imposing unrealistic parametric assumptions van der Laan (2015).
At the heart of HAL is the minimization of an empirical loss function, subject to a constraint on the function’s sectional variation norm, which effectively controls the model’s complexity. This combination allows HAL to achieve near-optimal rates of convergence and provides robust estimation of complex relationships between variables.
@@ 2.3.1 Representation of $Q$ in HAL: Zero-Order Approximation
where $P_{n}L(Q)$ is the empirical loss and $D^{(0)}_{M}([0,1]^{d})$ is the space of zero-order càdlàg functions with a bounded sectional variation norm $M$. This formulation ensures that the estimated function adheres to the constraint $||Q||^{*}_{v}<M$, promoting smoothness and preventing overfitting.
@@ 2.3.2 Approximation in Practice: Basis Expansion with Knot Points
In practice, HAL approximates $Q$ by constructing a large set of zero-order spline basis functions using knot points from the observed data. Given the input vector $X=(W,A)$, the basis functions are derived from the support points $\{X_{i}(S):1,\dots,n,S\subset\{1,\dots,d\}\}$. The resulting approximation is expressed as:
$\begin{split}Q_{\beta}(x)&=\beta_{0}+\sum_{s_{1}\subset\{1,2,...,d\}}\sum_{j=1}^{n}\beta_{s_{1},j}\phi_{u_{s},j}(x)=\boldsymbol{\Phi}^{0}(x)\boldsymbol{\beta}\\ \hskip 22.76228pt&\mbox{with}\hskip 22.76228pt||{\beta}||_{1}=|\beta_{0}|+\sum_{s_{1}\subset\{1,2,...,d\}}\sum_{j=1}^{n}|\beta_{s,j}|<M.\end{split}$
This approximation is a linear combination of zero-order basis functions $x\rightarrow\phi_{u_{s},j}(x)=I(x_{u_{s},j}\leq x_{s})$, defined by support points $x_{u_{s},j}$ from observations, along with their corresponding coefficient $\beta_{s,j}$.
@@ 2.3.3 Higher-Order HAL: First-Order and Higher
The first-order HAL approximates $Q$ using first-order spline basis functions, which allow for smoother representations of the target function. The general form of the first-order HAL approximation is:
@@
Extending beyond the first order, HAL can be generalized to accommodate $k$-th order smoothness classes $D^{(k)}([0,1]^{d})$, where $k=1,2,\ldots$. Functions in $D^{(k)}([0,1]^{d})$ can be approximated by a finite linear combination of $k$-th order splines. The $L_{1}$-norm corresponds with the sum over all $k$-th order derivatives of its sectional variation norms, referred to as the $k$-th order sectional variation norm. See works by van der Laan et al. (2022) for detailed theoretical development.
@@ 2.4.1 Sectional Variation Norm Bound (Penalty) Selector
The HAL estimator is constructed to minimize the empirical loss while selecting an appropriate bound for the sectional variation norm. The choice of the $L_{1}$-norm bound, $M$, is crucial for balancing bias and variance in the estimation. Cross-validation is used to select the optimal $M$. However, when estimating the dose-response curve, this often results in oversmoothed estimators with too much bias. Therefore, an undersmoothing strategy is employed, in which the bound is increased beyond the cross-validated value to reduce bias van der Laan et al. (2022).
In undersmoothed HAL, the $L_{1}$-norm bound $M$ is selected to be sufficiently large, larger than the value determined through cross-validated, to ensure the efficient influence curve score equation is solved at the level $\hat{\sigma}_{n}/(n^{1/2}\log n)$, where $\hat{\sigma}_{n}/n^{1/2}$ is the standard error estimate based on the sample variance of the efficient influence curve. This approach is analogous to selecting a smaller penalty, $\lambda$, in GLMNET, compared to the cross-validated value, while ensuring that the efficient influence curve criteria are met.
@@ 2.4.2 Smoothness Order Selector
In practical scenarios, the exact shape of the dose-response curve often unknown, posing challenges in determining the optimal level of smoothness and other crucial hyperparameters for the Highly Adaptive Lasso (HAL) fit, which depend on the curve’s characteristics. To address these challenges, the Smoothness-Adaptive HAL, introduced by van der Laan (2023), leverages the discrete Super Learner to identify the appropriate smoothness order and other relevant hyperparameters, such as the number of knots, within a cross-validation framework. This approach offers a robust data-driven method for selecting the hyperparameter that yield the best empirical performance in modeling the causal dose-response curve.

### v4

# HAL-Based Plug-in Estimation with Pointwise Asymptotic Normality of the Causal Dose-Response Curve

@@ Abstract
Estimating and obtaining reliable inference for the marginally adjusted causal dose-response curve for continuous treatments without relying on parametric assumptions is a well-known statistical challenge. Parametric models risk introducing significant bias through model misspecification, compromising the accurate representation of the underlying data and dose-response relationship. On the other hand, nonparametric models face difficulties as the dose-response curve is not pathwise differentiable, preventing consistent estimation at standard rates. The Highly Adaptive Lasso (HAL) maximum likelihood estimator offers a promising approach to this issue. In this paper, we introduce a HAL-based plug-in estimator for the causal dose-response curve, bridge theoretical development and empirical application, and assess its empirical performance against other estimators. This work emphasizes not just theoretical proofs, but also demonstrates their application through comprehensive simulations, thereby filling an essential gap between theory and practice. Our comprehensive simulations demonstrate that the HAL-based estimator achieves pointwise asymptotic normality with valid inference and consistently outperforms existing approaches for estimating the causal dose-response curve.
@@
A marginally adjusted dose-response curve, often referred to as a "causal" dose-response curve, provides a graphical representation of the population-level probability of a health outcome following a uniform application of a specific dose or, more generally, a dosage rule applied to the target population. Estimation of such a curve where the dose is ordinal or continuous requires accounting for potential confounders, missingness, and other typical complications in data-generating processes. Efficient estimators with valid inference are crucial for a wide range of scientific applications, including epidemiology, medical research, social sciences, and many other fields. For example, in a clinical trial investigating the impact of a continuous dose of drugs on blood pressure reduction, the dose-response curve can provide valuable insights into the specific reduction or increase in blood pressure attributed to each incremental change in dosage, accounting for confounding factors such as demographic and baseline clinical characteristics.
@@
To estimate the causal dose-response curve, we need to define a parameter under the counterfactual or potential outcome framework in which the outcome would be observed for subjects in the target population if all of them received a specific treatment level (van der Laan and Petersen, 2028; Petersen et al., 2006; Rubin, 2004). This framework, widely used in causal inference, has predominantly focused on binary interventions (van der Laan et al., 2011; Rubin, 2006; van der Laan and Robins, 2003; Abadie and Imbens, 2006; Angrist and Imbens, 1995; Robins, 1986; Rubin, 1978). Common approaches for estimating causal effects, such as the average treatment effect (ATE), include matching (Abadie and Imbens, 2011), estimating equations (e.g., inverse probability of treatment weighting (IPTW) (Austin and Stuart, 2015) and augmented IPTW (A-IPTW) (Luque-Fernandez et al., 2018)), and substitution estimators (e.g., G-computation (Robins, 1986) and targeted maximum likelihood estimation (TMLE) (van der Laan et al., 2011)). Furthermore, highly adaptive lasso (HAL) (van der Laan, 2017) and undersmoothed HAL (van der Laan et al., 2022) have been developed to provide reliable inference for the effects of binary treatment. Although significant progress has been made in the binary treatment context, many causal inference questions inherently involve continuous or ordinal interventions, such as doses in clinical treatments, for which consistent and efficient estimators remain scarce.
@@
The primary challenge in estimating the causal dose-response curve for continuous treatments in a nonparametric model is that the parameter is not pathwise differentiable (Bickel et al., 1993). Consequently, $\sqrt{n}$ estimators are not applicable, leading to reliance on estimators that depend on estimating densities rather than a smooth function of this density.
@@
Considerable efforts have been dedicated to crafting estimators for the causal dose-response curve. The Marginal Structural Model (MSM) framework, originally introduced by Robins et al. (2000), is known for defining causal dose-response curves in parametric models. However, these methods rely on specifying parametric models of the dose-reponse curve (MSM) and thus is subject to incorrect model specification. To address potential model misspecification, enhancements to MSM have been proposed using outcome prediction models, data-adaptive estimation of MSM, and projecting the true causal dose-response curve onto a functional space defined by MSM (Robins and Greenland, 2000; Haight et al., 2010; Neugebauer and van der Laan, 2007). Despite these advances, challenges related to robustness remain. Alternative estimators based on plug-in methods for conditional outcome models have been proposed (Hirano and Imbens, 2004; Callaway et al., 2021; Schwab et al., 2020; Díaz and van der Laan, 2013), but these often struggle with dimensionality issues, restrictive assumptions, and the complexity of valid inference. Techniques such as cross-validated TMLE (Chapter 25 in van der Laan and Rose, 2018) and nonparametric causal inference (Kennedy et al., 2017) achieve asymptotic normality, facilitating inference, but are constrained by their reliance on smoothness assumptions. Additionally, their reliance on kernel smoothing introduces complexities in selecting the tuning parameter, particularly given the suboptimal performance of global cross-validation selectors for smoothness and bandwidth selection (Díaz et al., 2023).
@@
In this article, we present and enhance the highly adaptive lasso (HAL) based plug-in estimator for the marginally adjusted causal dose-response, with a focus on bridging the gap between theory and empirical application. HAL-MLE is a nonparametric minimum-loss estimator capable of efficiently estimating pathwise differentiable parameters. Importantly, it has been theoretically proven to perform well with pointwise asymptotic normality in settings where the parameter is non-pathwise differentiable (van der Laan, 2023), such as with continuous treatments. Specifically, the HAL-based plug-in estimator converges to the true curve at a dimension-free rate up to log n factors (van der Laan, 2023). This work aims to implement the HAL-based estimator and validate its theoretical properties through empirical analysis.
@@
Through simulation studies, we assess the performance of this HAL-based estimator under various realistic conditions, showing that it provides valid inference with robust confidence intervals for the causal dose-response curve. These simulations confirm that the theoretical properties of the HAL estimator are effectively translated into practice, offering a thorough empirical validation of its utility for continuous dose-response estimation. Our contribution extends the utility of HAL to continuous dose-response problems, which have traditionally been difficult to estimate efficiently due to nonpathwise differentiability. We present the theoretical foundations of the HAL-based plug-in estimator and its associated delta-method-based inference that we implemented, followed by a series of simulations demonstrating its consistent performance and reliability.
@@
In the sections that follow, we define the target parameter and present the HAL-based plug-in estimator with confidence intervals based on the delta method. Then, we present simulation results that highlight the estimator’s asymptotic normality, nominal coverage, and overall reliability for estimating the marginally adjusted dose-response curve in continuous treatment settings.
@@ 1.1 Target Parameter and Statistical Model
Consider an observational study or a randomized experiment with $n$ independent and identically distributed observations of a continuous exposure or treatment variable $A$, an outcome $Y$, and covariates $W$, represented as $O=(W,A,Y)$. $O$ follows a true data-generating distribution (DGD) $P_{0}\in{\cal M}^{NP}$, where $\mathcal{M}^{NP}$ is a nonparametric model that places minimal assumptions on the data-generating process, thus allowing for a more flexible and realistic representation of complex relationships among variables. The goal is to estimate the marginal causal dose-response curve, which captures the causal relationship between treatment and outcome. The target parameter $\Psi_{P}(a)$ is defined as:
@@
When $Y$ is continuous, we focus on estimating the conditional mean $Q_{P}(W,A):=\mathbb{E}_{P}[Y|W,A]$, such that $\Psi_{P}(a)=\int_{w}Q(w,a)dP(w)$. This leads to the plug-in estimator $\hat{\Psi}(a)=\frac{1}{n}\sum_{i=1}^{n}\hat{Q}(W_{i},a)$, where $\hat{Q}$ is a nonparametric estimator of the conditional mean, such as the Highly Adaptive Lasso (HAL). When $Y$ is binary, we parameterize $E_{P}(Y\mid W,A)=1/(1+\exp(-Q_{P}(W,A))$ so that $Q_{P}$ now plays the role of $\mbox{Logit}E_{P}(Y\mid W,A)$ as the conditional probability.
Estimating this target parameter is challenging, since there is no efficient estimation method with reliable inference. Therefore, we adopt a flexible and realistic statistical model based on càdlàg functions with finite sectional variation norm. This approach allows for robust estimation without making unrealistic assumptions.
@@ 1.2 Flexible and Realistic Statistical Model: Càdlàg Functions
To model the relationship between $A$, $Y$, and $W$ without relying on strong parametric assumptions, we use the class of multivariate càdlàg (right-continuous with left limits) functions with finite sectional variation norms (FSVN) as the foundation of the statistical model. This class includes not only continuous functions but also functions with jumps and other non-linear features, providing high flexibility for approximating complex relationships in real-world data while avoiding strict smoothness assumptions.
@@
$Q||^{*}_{v}=|Q(0)|+\sum_{S\subset\{1,\dots,d\}}\int_{(0(S),1(S)]}|dQ_{S}(u)|.$
Càdlàg functions form a Donsker class, ensuring that empirical processes indexed by this class converge to Gaussian processes. This convergence provides desirable statistical properties, such as optimal convergence rates and asymptotic normality (Van der Vaart et al., 1996). These properties make càdlàg functions an attractive choice for constructing estimators in complex, high-dimensional data settings. In particular, Highly Adaptive Lasso (HAL) leverages this flexibility to estimate the target function $Q$ robustly, while achieving optimal convergence rates under weak assumptions.
@@ 1.3 Highly Adaptive Lasso (HAL) and Representation of $Q$: From Statistical Model to Practical Estimator
Building on the flexibility of càdlàg functions, the Highly Adaptive Lasso (HAL) is a nonparametric minimum-loss estimator that can flexibly estimate complex, high-dimensional target parameters. In its most nonparametric form, HAL only assumes that the true target function belongs to the class of $d$-variate real-valued càdlàg functions with bounded sectional variation norms. This allows HAL to fit a wide range of functions, including those with discontinuities and nonlinear relationships, making it a powerful tool for data-driven estimation without imposing unrealistic parametric assumptions (van der Laan, 2015).
At the heart of HAL is the minimization of an empirical loss function, subject to a constraint on the function’s sectional variation norm, which effectively controls the complexity of the model. This combination allows HAL to achieve near-optimal convergence rates and provides a robust estimation of complex relationships between variables.
@@ 1.3.1 Representation of $Q$ in HAL: Zero-Order Approximation
where $P_{n}L(Q)$ is the empirical loss and $D^{(0)}_{M}([0,1]^{d})$ is the space of zero-order càdlàg functions with a bounded sectional variation norm $M$. This formulation ensures that the estimated function adheres to the constraint $Q||^{*}_{v}<M$, promoting smoothness and preventing overfitting.
@@ 1.3.2 Approximation in Practice: Basis Expansion with Knot Points
In practice, HAL approximates $Q$ by constructing a large set of zero-order spline basis functions using knot points from the observed data. Given the input vector $X=(W,A)$, the basis functions are derived from the support points $\{X_{i}(S):1,\dots,n,S\subset\{1,\dots,d\}\}$. The resulting approximation is expressed as follows:
[…]
This approximation is a linear combination of zero-order basis functions $x\rightarrow\phi_{u_{s},j}(x)=I(x_{u_{s},j}\leq x_{s})$, defined by support points $x_{u_{s},j}$ from observations, along with their corresponding coefficients $\beta_{s,j}$.
@@ 1.3.3 Higher-Order HAL: First-Order and Higher
The first-order HAL approximates $Q$ using first-order spline basis functions, which allow smoother representations of the target function. The general form of the first-order HAL approximation is:
@@
Extending beyond the first order, HAL can be generalized to accommodate $k$-th order smoothness classes $D^{(k)}([0,1]^{d})$, where $k=1,2,\ldots$. Functions in $D^{(k)}([0,1]^{d})$ can be approximated by a finite linear combination of $k$-th order splines. The $L_{1}$-norm corresponds with the sum over all $k$-th-order derivatives of its sectional variation norms, referred to as the $k$-th order sectional variation norm. See works by van der Laan (2023) for detailed theoretical development.
@@ 1.4.1 Sectional Variation Norm Bound (Penalty) Selector
The HAL estimator is constructed to minimize the empirical loss while selecting an appropriate bound for the sectional variation norm. The choice of the $L_{1}$-norm bound, $M$, is crucial for balancing bias and variance in the estimation. Cross-validation is used to select the optimal $M$. However, when estimating the dose-response curve, this often results in oversmooth estimators, biased for the functional of interest. Therefore, an undersmoothing strategy is employed, in which the bound is increased beyond the cross-validated value to reduce bias (van der Laan et al., 2022).
In undersmoothed HAL, the $L_{1}$-norm bound $M$ is selected to be sufficiently large, larger than the value determined through cross-validated, to ensure the efficient influence curve score equation is solved at the level $\hat{\sigma}_{n}/(n^{1/2}\log n)$, where $\hat{\sigma}_{n}/n^{1/2}$ is the standard error estimate based on the sample variance of the efficient influence curve. This approach is analogous to selecting a smaller penalty, $\lambda$, in GLMNET, compared to the cross-validated value, with the goal of solving the efficient influence curve score equations.
@@ 1.4.2 Smoothness Order Selector
In practical scenarios, the exact shape of the dose-response curve is often unknown, posing challenges in determining the optimal level of smoothness and other crucial hyperparameters for the Highly Adaptive Lasso (HAL) fit, which depend on the curve’s characteristics. To address these challenges, the Smoothness-Adaptive HAL, introduced by van der Laan (2023), leverages the discrete Super Learner to identify the appropriate smoothness order and other relevant hyperparameters, such as the number of knots, within a cross-validation framework. This approach offers a robust data-driven method for selecting the hyperparameter that yields the best empirical performance in modeling the causal dose-response curve.